### v1

# MAPo : Motion-Aware Partitioning of Deformable 3D Gaussian Splatting for High-Fidelity Dynamic Scene Reconstruction

Abstract
3D Gaussian Splatting, known for enabling high-quality static scene reconstruction with fast rendering, is increasingly being applied to dynamic scene reconstruction. A common strategy involves learning a deformation field to model the temporal changes of a canonical set of 3D Gaussians. However, these deformation-based methods often produce blurred renderings and lose fine motion details in highly dynamic regions due to the inherent limitations of a single, unified model in representing diverse motion patterns. To address these challenges, we introduce Motion-Aware Partitioning of Deformable 3D Gaussian Splatting (MAPo), a novel framework for high-fidelity dynamic scene reconstruction. Its core is a dynamic score-based partitioning strategy that distinguishes between high- and low-dynamic 3D Gaussians. For high-dynamic 3D Gaussians, we recursively partition them temporally and duplicate their deformation networks for each new temporal segment, enabling specialized modeling to capture intricate motion details. Concurrently, low-dynamic 3DGs are treated as static to reduce computational costs. However, this temporal partitioning strategy for high-dynamic 3DGs can introduce visual discontinuities across frames at the partition boundaries. To address this, we introduce a cross-frame consistency loss, which not only ensures visual continuity but also further enhances rendering quality. Extensive experiments demonstrate that MAPo achieves superior rendering quality compared to baselines while maintaining comparable computational costs, particularly in regions with complex or rapid motions.
Figure 1: Overview. (a-b) Deformation-based methods often blur details in regions with complex or rapid motion. (c) Our MAPo significantly improves rendering quality in these areas. (d) Ground Truth.
1 Introduction
Dynamic scene reconstruction is a fundamental challenge in computer vision, with broad applications in virtual reality, visual effects, and autonomous driving, among other applications. In recent years, Neural Radiance Fields (NeRF) (Mildenhall et al. 2021) have demonstrated remarkable capabilities in representing static scenes for novel view synthesis by leveraging implicit neural representations. Building upon this, numerous efforts (Pumarola et al. 2021; Xian et al. 2021; Park et al. 2021b, a; Li et al. 2022a, b; Fang et al. 2022; Fridovich-Keil et al. 2023; Cao and Johnson 2023; Shao et al. 2023) have extended NeRF to dynamic scenes by introducing temporal conditioning. However, the inherent reliance on dense spatial sampling and costly Multilayer Perceptron (MLP) querying leads to significant limitations in both training efficiency and rendering speed.
Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) has emerged as a powerful alternative for static scene reconstruction, achieving real-time rendering while maintaining photorealistic rendering quality. This method explicitly represents scenes using anisotropic 3D Gaussian (3DG) primitives. Encouraged by its success in static scenes, researchers have begun adapting Gaussian Splatting to dynamic scenes (Yang et al. 2023; Wu et al. 2023; Huang et al. 2023; Bae et al. 2024; Wu et al. 2025; Lu et al. 2024; Lee et al. 2024; Yang et al. 2024). Employing deformable 3D Gaussians—where a learned deformation field maps a canonical set of 3D Gaussians(3DGs) to their time-varying states—has quickly emerged as a prevailing and effective strategy for reconstructing dynamic scenes, offering compelling visual quality from a compact representation. Despite promising results, these methods suffer from two critical limitations inherent in their deformation framework:
  • Bottleneck in Motion Modeling Capacity: As shown in Fig. 1, deformation-based methods often produce blurry reconstructions and lose fine motion details in regions with complex or rapid motions. This core limitation stems from their unified modeling strategy, which relies on a single canonical set of 3DGs and a single, globally shared deformation network to represent all spatio-temporal variations. Such an approach forces the network to find a single set of parameters that best fits all diverse and often conflicting motion patterns across the entire time sequence. Consequently, the learned representation often converges to a temporally averaged representation. As shown in Fig. 2, this averaging effect prevents the model from accurately capturing complex motions that deviate from this learned average, such as abrupt changes and fine-grained temporal details. This ultimately leads to the blurring and loss of fine details in dynamic regions.
  • Redundant Computation: The 3DGs in static regions still participate in repeated deformation network computation, wasting computational resources and potentially slowing down training and rendering speed.
To tackle these issues, we introduce MAPo, a novel framework for high-fidelity dynamic scene reconstruction. Its core is a dynamic score-based partitioning strategy comprising two key components: Temporal Partitioning Based on Dynamic Scores and Static 3D Gaussian Partitioning. The former enhances details in dynamic regions through hierarchical temporal partitioning of high-dynamic 3DGs, and the latter resolves redundant computation by treating low-dynamic 3DGs as static. To ensure temporal smoothness after partitioning, a cross-frame consistency loss is also employed.
Specifically, MAPo first analyzes each 3DG’s historical positions to compute a dynamic score based on its maximum displacement and variance, providing a practical measure of motion intensity. During temporal partitioning, high-dynamic 3DGs, identified by their high dynamic scores, are recursively partitioned along the temporal dimension. The 3DGs that undergo partitioning are replicated, creating new 3DG instances for the resulting temporal sub-segments. Simultaneously, the deformation network of the parent segment is replicated, creating a dedicated sub-network for each new temporal sub-segment to deform all 3DGs within it. Instead of relying on a single, unified model, our strategy enables multiple sets of networks and their corresponding 3DGs to model the dynamic scene over distinct temporal segments. This allows for a more faithful capture of details in highly dynamic regions and effectively alleviates the “temporal averaging” effect. Meanwhile, 3DGs with low dynamic scores are identified as static. Their attributes are directly updated to reflect the deformation results, after which redundant deformation network computations are skipped to reduce computational costs.
To further mitigate visual discontinuities introduced by temporal partitioning, we introduce a cross-frame consistency loss that enforces two constraints: (i) the renderings of two sets of temporally adjacent 3DGs should remain as consistent as possible when rendered at the same time and from the same viewpoint, and (ii) the rendering of the temporally adjacent 3DGs corresponding to the current time should also align with the ground-truth observation at the current time. This consistency loss not only reduces visual discontinuities at partition boundaries but also leverages temporal context for more faithful reconstructions. Our key contributions are summarized as follows:
  • We propose MAPo, a novel framework for high-fidelity dynamic scene reconstruction based on a dynamic score-based partitioning strategy. Our strategy enhances the modeling effects of highly dynamic regions by enabling multiple sets of networks and their corresponding 3DGs to model the dynamic scene over distinct temporal segments, and simultaneously reduces computational costs by identifying low dynamic scores as static.
  • We design a cross-frame consistency loss that effectively mitigates visual discontinuities introduced by temporal partitioning and further enhances the rendering quality.
  • Extensive experiments demonstrate that MAPo achieves state-of-the-art (SOTA) rendering quality, particularly in capturing fine details in highly dynamic scenes.
(a) Average
(b) Rendering 1
(c) Rendering 2
Figure 2: Rendering Results of a Single Unified Model. (a) shows the temporally averaged representation, which is visualized by directly rendering the canonical 3DGs without any deformation. The regions highlighted in blue in (b) and (c) are close to this average. The region highlighted in red in (c) is distant from this average.
2 Related Work
Dynamic NeRF
Neural Radiance Fields (NeRF) (Mildenhall et al. 2021) marked a significant breakthrough in representing static scenes for high-quality novel view synthesis using implicit MLPs. This success has spurred numerous efforts to extend its capabilities to dynamic scenes. One prominent strategy is deformation-based, where a learned deformation field maps observed points to a canonical representation. D-NeRF (Pumarola et al. 2021) pioneered this concept by directly using spatio-temporal coordinates as the input to the deformation network. This was later advanced by methods like Nerfies (Park et al. 2021a), replacing the explicit time input with a learnable latent code to govern deformation, and HyperNeRF (Park et al. 2021b), utilizing higher-dimensional embeddings to better model complex topologies. NeRFPlayer (Song et al. 2023) introduces a streaming-capable representation that separates static and dynamic scene components. Other approaches bypass the deformation-based paradigm. For instance, DyNeRF (Li et al. 2022b) directly models the scene’s temporal evolution by conditioning a 6D plenoptic function on a temporal latent code. To improve efficiency, a different line of work, including K-Planes (Fridovich-Keil et al. 2023) and HexPlane (Cao and Johnson 2023), leverages explicit feature representations like multi-planar decompositions. HyperReel (Attal et al. 2023) builds on this principle with a compact dual-plane representation to further enhance rendering speed. Despite these advances, the reliance on dense ray sampling and per-point MLP queries in NeRF-based methods remains a barrier to achieving real-time rendering.
Dynamic Gaussian Splatting
The recent success of 3DGS in achieving real-time, photorealistic rendering for static scenes has naturally inspired its extension to dynamic scenes. Among these extensions, a significant number of approaches are deformation-based, where a canonical set of 3DGs is deformed over time. D3DGS (Yang et al. 2023) introduces a canonical representation and a deformation network driven by position and time, to map canonical 3DGs to the observation space. Many subsequent methods share similarities with this paradigm. 4DGaussians (Wu et al. 2023) decodes features from a HexPlane representation using spatio-temporal coordinates to derive per-Gaussian deformations. E-D3DGS (Bae et al. 2024) advances this by replacing direct coordinate inputs with temporal and per-Gaussian embeddings and adopting a dual-deformation strategy. Similarly, DN-4DGS (Lu et al. 2024) integrates spatio-temporal information from neighboring elements and also employs a dual-deformation approach. To handle long sequences, SWinGS (Shaw et al. 2024) partitions the sequence into sliding windows and trains an independent model for each, using 2D optical flow to guide window segmentation and a post-hoc fine-tuning step to ensure smoothness. To improve efficiency, Swift4D (Wu et al. 2025) uses 2D RGB information to separate dynamic from static regions, modeling only the dynamic 3DGs’ deformation with a 4D hash grid and a small MLP decoder. However, these deformation-based methods, despite their diversity, face a fundamental trade-off. Approaches relying on a single, globally shared representation like D3DGS and its direct successors often suffer from the “temporal averaging” effect in complex scenes. Conversely, strategies like SWinGS that partition the sequence into independent, localized models mitigate this issue, but introduce their own challenges, including coarse, window-level partitioning; cumbersome pre- and post-processing steps; and a strong reliance on 2D priors like optical flow.
Other distinct strategies have also been explored. Curve-based methods (Gao et al. 2024b; Li et al. 2023; Lee et al. 2024) model 3DG attribute changes over time using parametric curves, but may induce drift under complex motions and introduce significant storage overhead. 4D Gaussian-based method 4DGS (Duan et al. 2024) decomposes 4D Gaussians into 3DGs and marginal 1D Gaussians, offering a direct spatio-temporal representation, but incurs significant computational costs. Per-frame training methods (Sun et al. 2024; Gao et al. 2024a; Hu et al. 2025) enable online reconstruction but are constrained by the inherent trade-off between quality, storage, and efficiency.
To overcome the limitations of existing approaches, we introduce a dynamic score-based partitioning strategy. Instead of relying on a single, unified model, we recursively subdivide high-dynamic 3DGs into finer temporal segments to address the issue of “temporal average”. Unlike window-based methods like SWinGS, our partitioning operates on a fine-grained, per-3DG level, is guided directly by 3D motion cues instead of 2D priors, and is integrated into a single, end-to-end training framework, thus avoiding cumbersome pre- and post-processing steps. Furthermore, MAPo is inherently built upon a compact deformation field architecture, and it identifies and treats low-dynamic 3DGs as static to avoid the substantial overhead of other categories of methods.
3 Preliminaries
3D Gaussian Splatting
3D Gaussian Splatting introduces an explicit point-based representation where each point in the 3D point cloud is equipped with four fundamental properties: mean $\mu$, covariance matrix $\Sigma$, opacity $\alpha$, and spherical harmonics $sh$. The covariance matrix $\Sigma$ can be decomposed as $\Sigma=RSS^{T}R^{T}$, where $S$ is parameterized by a 3D vector $s$ and $R$ is parameterized by a quaternion $q$. The rendering process is fully differentiable. For a given viewpoint, all 3DGs are first projected onto the 2D image plane. Then, these projected 2D Gaussians are sorted by depth and blended together using alpha-compositing to synthesize the final pixel color $C$:
$C=\sum_{i=1}^{N}c_{i}\alpha_{i}^{\prime}\prod_{j=1}^{i-1}\left(1-\alpha_{j}^{\prime}\right).$ (1)
where $c_{i}$ is the view-dependent color computed from $sh$ and viewpoint, and $\alpha^{\prime}_{i}$ is the 2D evaluation of the 3DG’s opacity.
Embedding-Based Deformation
Our work builds upon the coarse-to-fine deformation mechanism introduced in E-D3DGS (Bae et al. 2024). In this paradigm, each 3DG is associated with a learnable embedding $\mathbf{z}_{g}$, and each timestamp $t$ is represented by a pair of temporal embeddings: a coarse embedding $\mathbf{z}_{t_{c}}$ capturing low-frequency motion, and a fine embedding $\mathbf{z}_{t_{f}}$ for high-frequency details. A coarse deformation network $\mathcal{F}$ and a fine deformation network $\mathcal{F}_{\theta}$ process these embeddings to predict deformations for all 3DG attributes. The final deformation is the sum of the coarse and fine predictions:
$(\Delta{\mu},\Delta{q},\Delta{s},\Delta{\alpha},\Delta{sh})=\mathcal{F}(\mathbf{z}_{g},\mathbf{z}_{t_{c}})+\mathcal{F}_{\theta}(\mathbf{z}_{g},\mathbf{z}_{t_{f}}).$ (2)
Figure 3: An overview of MAPo. (a) 3DGs’ deformation process. (b) Compute the dynamic score of 3DGs from history positions during training. (c) High-dynamic 3DGs are recursively temporally partitioned, and low-dynamic ones are deformed and treated as static. (d) Dynamic and static 3DGs are combined for rendering. Losses are computed on the left.
4 Method
Our approach consists of two main components: a dynamic score-based partitioning strategy and a cross-frame consistency loss. The overview of our method is shown in Fig. 3. First, we elaborate on the Dynamic Score-based Partitioning Strategy, including how we compute a dynamic score for each 3DG based on its historical positions, how this score guides a partitioning where high-dynamic 3DGs are recursively temporally partitioned, and how low-dynamic 3DGs are identified as static. Subsequently, we describe our Cross-Frame Consistency Loss, which is designed to address the visual discontinuities caused by partitioning.
Dynamic Score-based Partitioning Strategy
Dynamic Score Calculation
To accurately characterize the motion intensity of each 3DG across temporal variations, we design a comprehensive scoring mechanism that integrates both maximum displacement and position variance. For each 3DG $G_{i}$, we record its spatial position $\mu_{ij}$ during training, where $i\in\{1,2,\cdots,N\}$ denotes the index of 3DG and $j\in\{1,2,\cdots,m\}$ represents the index of the recorded historical positions. Here, $m$ is a hyperparameter controlling the number of recorded positions per 3DG. Initially, we considered using the maximum displacement as the sole metric for quantifying the motion intensity of 3DGs, which we efficiently compute as the length of the diagonal of the historical positions’ axis-aligned bounding box. However, upon closer examination of the positional records, we realized that relying solely on this metric would be insufficient. For instance, an object with short-term high-speed motion but long-term stillness may exhibit a large maximum displacement but relatively low variance. Conversely, an object with continuous small oscillations would have a small maximum displacement but a relatively high variance, suggesting a consistent and complex yet less pronounced movement pattern. Therefore, to comprehensively characterize diverse motion behaviors to accurately identify more challenging motion, we employ two quantitative metrics for each 3DG. For the i-th 3DG, its maximum displacement $r_{i}$ and position variance $v_{i}$, computed as:
$r_{i}=\bigl{\|}\max_{j}\mu_{ij}-\min_{j}\mu_{ij}\bigr{\|},\quad v_{i}=\sum_{j=1}^{m}\frac{\|\mu_{ij}-\bar{\mu}_{i}\|^{2}}{m},$ (3)
Here, $\max_{j}\mu_{ij}$ and $\min_{j}\mu_{ij}$ are the element-wise maximum and minimum vectors over the historical positions $\{\mu_{ij}\}_{j=1}^{m}$ for 3DG $i$. The term $\bar{\mu}_{i}=\frac{1}{m}\sum_{j=1}^{m}\mu_{ij}$ represents the mean position. The maximum displacement $r_{i}$ captures the peak amplitude of motion, while the variance $v_{i}$ measures the dispersion around the mean position. After computing the maximum displacement and variance, we map all 3DGs’ $r$ and $v$ values to the interval $[0,1]$ via percentile-based normalization:
$\tilde{r}_{i}=\sum_{k=1}^{100}\frac{\mathbf{1}(r_{i}\geq q_{r}(k))}{100},\quad\tilde{v}_{i}=\sum_{k=1}^{100}\frac{\mathbf{1}(v_{i}\geq q_{v}(k))}{100},$ (4)
where $\mathbf{1}(\cdot)$ denotes the indicator function, and $q_{r}(k)$ and $q_{v}(k)$ are the $k$-th percentiles of $\{r_{i}\}$ and $\{v_{i}\}$, respectively. We use the harmonic mean to fuse $\tilde{r}_{i}$ and $\tilde{v}_{i}$, as it requires both inputs to be high for a high output. The final dynamic score $S_{i}$ of the i-th 3DG is denoted as:
$S_{i}=\frac{2}{\frac{1}{\tilde{r}_{i}+\varepsilon}+\frac{1}{\tilde{v}_{i}+\varepsilon}},$ (5)
where $\varepsilon=10^{-6}$ is used for numerical stability.
Temporal Partitioning Based on Dynamic Scores
3DGs with high dynamic scores typically correspond to regions with complex or rapid motion. Since a single 3DG struggles to effectively model long-term or complex motion characteristics, an intuitive and effective solution is to partition these high-dynamic 3DGs along the temporal dimension based on their dynamic scores, thereby capturing motion variations more precisely. Fig. 4 demonstrates the effectiveness of this partitioning through a simple demo.
Figure 4: Effectiveness of Temporal Partitioning Strategy and Consistency Loss on a Toy example. (a) A 3D curve $\mathbf{p}(t)$ simulates a dynamic trajectory. (b) A single point and a single MLP to fit $\mathbf{p}(t)$ for the entire duration; (c) Two points and two corresponding MLPs for two partitioned time segments; (d) Apply a consistency loss to (c) at the partition boundary.
Building on this insight, we recursively partition the 3DGs along the temporal dimension, progressively subdividing those with higher dynamic scores into finer temporal segments to achieve more precise dynamic reconstruction. For the complete time range $[0,T]$, each 3DG maintains two key properties: its partition level $l$ (initialized to 0) and its temporal segment range $t_{\text{start}},t_{\text{end}}]$ (initially $[0,T]$). For notational convenience, all such ranges are treated as left-closed and right-open. Let $G_{[t_{\text{start}},t_{\text{end}}]}$ denote the set of 3DGs assigned to the temporal segment range $t_{\text{start}},t_{\text{end}}]$. Let $F_{[t_{\text{start}},t_{\text{end}}]}$ represent the deformation network corresponding to the temporal segment range $t_{\text{start}},t_{\text{end}}]$. It is responsible for deforming the 3DGs in $G_{[t_{\text{start}},t_{\text{end}}]}$. When a 3DG at level $l$ exhibits a dynamic score exceeding the current-level threshold $\tau_{l}$ within its time range $t_{\text{start}},t_{\text{end}}]$, we partition it at the temporal midpoint $t_{\text{mid}}=(t_{\text{start}}+t_{\text{end}})/2$. The original 3DG retains the first sub-segment $t_{\text{start}},t_{\text{mid}}]$ while advancing to level $l+1$, and a new replica is created for the second sub-segment $t_{\text{mid}},t_{\text{end}}]$ with identical attributes. Correspondingly, the deformation network $F_{[t_{\text{start}},t_{\text{end}}]}$ is replicated to create $F_{[t_{\text{start}},t_{\text{mid}}]}$ and $F_{[t_{\text{mid}},t_{\text{end}}]}$ to model the distinct spatio-temporal deformation patterns within each sub-segment. Within each new shorter temporal segment, this partitioning process is applied recursively. Driven by the complexity of motion in different regions, our hierarchical approach adaptively refines their corresponding temporal segments.
Static 3D Gaussian Partitioning
3DGs with dynamic scores below a predefined threshold $\tau_{\text{static}}$ are identified as static. These static 3DGs have their attributes initialized once using the output of their associated deformation network at a randomly sampled timestep. Subsequently, they are excluded from computations involving the deformation network during rendering while their attributes remain optimizable, significantly reducing computational costs.
(a) 4DGaussians
(b) Ex4DGS
(c) Swift4D
(d) E-D3DGS
(e) Ours
(f) Ground Truth
Figure 5: Qualitative comparisons against existing SOTA methods on the MeetRoom and N3DV dataset.
Cross-Frame Consistency Loss
Temporal partitioning of 3DGs, while beneficial for modeling complex motions, can introduce visual discontinuities at the partition boundaries. To ensure temporal smoothness, we introduce the cross-frame consistency loss ${L}_{\text{cross}}$, which consists of two components: ${L}_{\text{current}}$ and ${L}_{\text{gt}}$. The ${L}_{\text{current}}$ evaluates rendering consistency at the partition boundary. It compares two renderings of the same frame at this boundary:
${L}_{\text{current}}=\left\|I_{t}(G_{t},V)-I_{t}(G_{t^{{}^{\prime}}},V)\right\|_{1}.$ (6)
Here, $G_{t}$ denotes the set of all 3DGs used for rendering at a timestamp $t$. $t^{\prime}$ is the timestamp from the nearest neighboring temporal segment closest to $t$. $G_{t}$ comprises two distinct subsets: static 3DGs and dynamic 3DGs. The dynamic 3DGs active at timestamp $t$, denoted as $G_{t}^{d}$, are governed by their corresponding deformation networks $F_{t}^{d}$. $I_{t}(G_{t},V)$ denotes the image rendered at timestamp $t$ from viewpoint $V$ using $G_{t}$. By computing the L1 norm between these two rendered images, ${L}_{\text{current}}$ captures the discrepancy between adjacent temporal segments in rendering the same frame. Optimizing this term helps reduce visual discontinuities, ensuring smoother transitions between temporal segments.
Although ${L}_{\text{current}}$ effectively mitigates visual discontinuities across partition boundaries, we observe that relying on it alone can compromise rendering quality. Since ${L}_{\text{current}}$ only enforces self-consistency between adjacent segments without an external reference, continuous optimization can cause them to converge to a consistent but over-smoothed state, leading to perceptible blurring in dynamic regions. To counteract this and anchor the reconstruction to the ground truth, we introduce ${L}_{\text{gt}}$. This loss is designed to enrich the adjacent segment’s 3DGs with valuable spatio-temporal context from the current frame. It achieves this by directly supervising their rendering against the ground-truth image of the current view $V$, which is captured at timestamp $t$:
${L}_{\text{gt}}=\left\|I_{t}(G_{t^{{}^{\prime}}},V)-I^{\text{GT}}\right\|_{1},$ (7)
where $I^{\text{GT}}$ denotes the corresponding ground-truth image at timestamp $t$. This contextual enrichment forces $G_{t^{\prime}}$ to learn to represent the sharp details of the current frame, thereby preventing over-smoothing and enhancing overall fidelity.
Finally, the overall cross-frame consistency loss, ${L}_{\text{cross}}$, is defined as a weighted combination of ${L}_{\text{current}}$ and ${L}_{\text{gt}}$:
${L}_{\text{cross}}=0.5\cdot{L}_{\text{current}}+{L}_{\text{gt}},$ (8)
By optimizing ${L}_{\text{cross}}$, we reduce discontinuities across temporal segments while enhancing rendering quality. We apply ${L}_{\text{cross}}$ only for training views whose frame indices are within $5$ frames of any partition boundary. Fig. 4 (d) demonstrates its effectiveness through a simple demo.
[…]
DyNeRF1,2 29.58 - 0.083 56MB 1344 hours 0.01
NeRFPlayer1,3 30.69 0.932 0.111 1654MB 5 hours 36 mins 0.06
Mix Voxels 30.30 0.918 0.127 512 MB 1 hour 28 mins 1.01
K-Planes 30.86 0.939 0.096 309 MB 1 hour 33 mins 0.15
HyperReel4 30.37 0.921 0.106 1362 MB 8 hours 42 mins 1.19
D3DGS 28.27 0.917 0.156 75MB 2 hours 17 mins 20.29
4DGS 30.30 0.933 0.069 3.6GB 7 hours 43 mins 54.36
4DGaussians 30.19 0.917 0.061 53 MB 1 hour 13 mins 78.28
Ex4DGS 30.76 0.939 0.056 205 MB 1 hour 5 mins 51.46
Swift4D 30.05 0.931 0.055 116MB 48 mins 138.00
4DGC 30.78 0.938 0.052 225 MB 5 hours 44 mins 124.61
E-D3DGS 30.79 0.934 0.051 73 MB 2 hours 41 mins 37.51
Ours 31.26 0.943 0.044 65 MB 1 hour 52 mins 75.64
Table 1: Quantitative comparison on the N3DV dataset. 1 flame_salmon was trained on only frag1. 2 only reported results on the flame_salmon_frag1 and was trained on 8 GPUs. 3 trained with 90 frames. 4 trained with 50 frames. Storage, training time, and FPS are measured on flame_salmon_frag1.
[…]
D3DGS 25.81 0.890 0.233 36 MB 47 mins 42.51
4DGS 26.12 0.896 0.080 5.4 GB 6 hours 32 mins 70.54
4DGaussians 26.16 0.894 0.081 51 MB 14 mins 31.98
Ex4DGS 26.46 0.895 0.083 123 MB 1 hour 6 mins 117.49
Swift4D 25.51 0.882 0.085 76 MB 20 mins 109.58
4DGC 26.56 0.901 0.070 224MB 3 hours 26 mins 160.59
E-D3DGS 26.24 0.896 0.081 28 MB 1 hour 36 mins 90.26
Ours 26.72 0.903 0.066 49 MB 1 hour 19 mins 92.21
Table 2: Quantitative comparison on MeetRoom dataset. Storage, training time, and FPS are calculated on discussion.
5 Experiment
Dataset and Metrics
We evaluate our method on two real-world dynamic scene datasets: N3DV (Li et al. 2022b) and Meet Room (Li et al. 2022a). The N3DV dataset includes videos (typically 10s, 30 FPS) captured by 20 cameras. For fair comparison, we downsample its images to 1352×1014 and segment the longer flame_salmon sequence into four 10s clips. The Meet Room dataset consists of 10s, 30 FPS videos (1280×720) from 13 cameras. We report PSNR, SSIM, and LPIPS for rendering quality and also assess computational costs, including training time, rendering speed, and storage, which is calculated on NVIDIA RTX A6000.
Implementation Details
Our implementation builds upon the E-D3DGS codebase. We provide comprehensive technical details in the appendix, which covers specific implementation details, hyperparameter settings, in-depth analyses of our method’s mechanics, and extensive supplementary experiments, including more baseline comparisons and additional ablation studies.
Comparisons
Quantitative Comparisons.
We compare MAPo against SOTA NeRF-based and 3DGS-based methods. To ensure fairness, all 3DGS-based baselines are evaluated under consistent conditions, including identical point cloud initializations. Direct comparisons with methods employing significantly different protocols (e.g., SWinGS, ST-GS) are deferred to the appendix, where full details on all baselines and experimental setups are also provided. As shown in Tab. 1 and Tab. 2, our method consistently achieves SOTA rendering quality across both datasets. We attribute this superior performance to our dynamic score-based partitioning strategy. Unlike approaches that rely on a single, unified deformation model (e.g., D3DGS, E-D3DGS), MAPo’s hierarchical partitioning enables specialized, independent modeling for high-dynamic 3DGs. This effectively mitigates the “temporal averaging” effect, allowing for the reconstruction of fine-grained details that are often blurred or lost in other methods, as evidenced by our top-ranking PSNR, SSIM, and LPIPS scores on both datasets. Furthermore, MAPo achieves this SOTA rendering quality while maintaining competitive computational efficiency. Our static 3D Gaussian partitioning significantly reduces overhead by eliminating redundant computations for low-dynamic 3DGs. As a result, MAPo offers a compelling balance between high fidelity and practical resource usage.
Qualitative Comparisons.
We present qualitative results in Fig. 5. The comparison highlights that baseline methods often produce degraded results in areas with complex motion. For example, in cases with fast-moving hands or detailed facial expressions, baseline methods exhibit severe motion blur and loss of detail. This is largely because their single, unified model struggles to accurately capture the diverse, complex motions. Benefiting from our temporal partitioning strategy, our approach preserves fine details in highly dynamic regions, yielding renderings that are significantly sharper and more faithful to the ground truth.
[…]
Table 3: Progressive Component Ablation on Vrheadset.
Baseline
(1.1) + Max Dis
(1.2) + Var
Ground Truth
Figure 6: Observation of Dynamic Partition on Vrheadset.
(1.2) + Var
(2.0) + Static
Static Region
Figure 7: Observation of Static Partition on Salmon.
Ablation Study and Analysis
We present the quantitative results of progressively incorporating each component of our method in Tab. 3. These components include: 1. Temporal Partition, referring to Temporal Partitioning Based on Dynamic Scores, where we initially employ Maximum Displacement (1.1 + Max Dis) for dynamic score computation to partition dynamic 3DGs, followed by incorporating Variance (1.2 + Var) for the full dynamic score; 2. Static Partition, denoting Static 3D Gaussian Partitioning; and 3. Cross, representing Cross-Frame Consistency Loss, which involves the sequential integration of ${L}_{\text{current}}$ (3.1 + ${L}_{\text{current}}$) and ${L}_{\text{gt}}$ (3.2 + ${L}_{\text{gt}}$). Each component is added incrementally, building upon the previously incorporated modules to evaluate their collective impact on performance. For clarity, in the subsequent experiments, we use consistent numbering to denote the variants at each stage.
Annotation
(2.0) + Static
(3.1) + ${L}_{\text{current}}$
(3.2) + ${L}_{\text{gt}}$
Ground Truth
Figure 8: Observation of ${L}_{\text{cross}}$’s Effect on Jitter. The four images on the right are obtained by vertically concatenating the horizontal line in the “Annotation” over time.
(2.0) + Static
(3.1) + ${L}_{\text{current}}$
(3.2) + ${L}_{\text{gt}}$
Ground Truth
Figure 9: Observation of ${L}_{\text{cross}}$’s Effect on Quality.
Temporal Partition. Tab. 3 and Fig. 6 clearly demonstrate the quality improvements achieved through temporal partition. By leveraging maximum displacement to compute dynamic scores for partitioning, we significantly enhance rendering quality. Further incorporating variance into the dynamic score computation yields better rendering quality.
Static Partition. As shown in Tab. 3, Static Partition reduces computational costs compared with Temporal Partition. Additionally, Fig. 7 shows that the visual quality of static regions in Static Partition remains comparable to that of Temporal Partition without static partition.
Cross. Temporal partitioning can lead to discontinuities between frames near partition boundaries, as illustrated in Fig. 8 (2.0), where the region highlighted by a red box reveals a sharp visual discontinuity that manifests as a horizontal seam. By introducing ${L}_{\text{current}}$ and ${L}_{\text{gt}}$, we mitigate this issue. Fig. 9 shows that our cross-frame consistency loss also enhances the rendering quality, reconstructing finer details like the spray gun’s barrel, which was absent in (2.0).
6 Conclusion
We propose a novel method for high-fidelity dynamic scene reconstruction. Based on deformation-based 3DGS, MAPo leverages a dynamic score computed from historical positions to guide a hierarchical partitioning strategy that addresses the limitations of unified models. In addition, we introduce a cross-frame consistency loss to effectively mitigate visual discontinuities introduced at partition boundaries. This adaptive partitioning approach enables specialized, high-fidelity reconstruction for complex motion regions while efficiently handling static parts of the scene. Extensive experiments demonstrate that MAPo achieves state-of-the-art rendering quality, particularly excelling at capturing intricate motion details in challenging dynamic scenes while maintaining comparable computational efficiency.